OpenStudy (anonymous):

The theme park company is building a scale model of the killer whale stadium main show tank for an investor's presentation. Each dimension will be made 6 times smaller to accommodate the mock-up in the presentation room. How many times smaller than the actual volume is the volume of the mock-up?

OpenStudy (anonymous):

Each tank is a half cylinder, Im so confused

OpenStudy (anonymous):

SF = AF^2 = VF^3

OpenStudy (anonymous):

I don't understand? I haven't learned anything like that

OpenStudy (anonymous):

That means scale factor (1d) = area factor (2d) ^2 = volume factor (3d) ^3

OpenStudy (anonymous):

Oh ok, so does that mean it will be 216 times smaller?

OpenStudy (anonymous):

So if each dimension, which is a 1d measurement is made 6 times smaller, so 1/6 of the size, the volume will have changed by that ^3'd. So 1/6x1/6x1/6 which = 1/216 of the volume

OpenStudy (anonymous):

How am I supposed to show what the percentage is?

OpenStudy (anonymous):

I think you just multiply by 100 to change a fraction to a percentage

OpenStudy (anonymous):

Can you show I m still confused??/

OpenStudy (anonymous):

I understand what you said before just not the percentage thing

OpenStudy (anonymous):

Let's take the nice easy fraction 3/10 as an example. We know this is 30% just because that's what it is. But if we multiply 0.3 by 100 we also get 30

OpenStudy (anonymous):

Ok so is it 216/10 = 26.1 so it is 26.1 percent the size of the original? Or smaller, 2.61 percent?

OpenStudy (anonymous):

Well it's 216 times smaller, so that is 1/216 of the original volume. Multiply that by 100 to get 0.4629% of the volume

OpenStudy (anonymous):

But what are you multiplying to 100.......

OpenStudy (anonymous):

1/216 * 100

OpenStudy (anonymous):

How are you getting 4629 from that

OpenStudy (anonymous):

100/216?

OpenStudy (anonymous):

Ok i get it now thanks for the help